# secReferences

### v1

@@ Appendix A Proofs
For any vertex type $(b_{1},b_{2},r,1)$ and $(b_{1},b_{2},r,0)$, the expected number of these vertices in $G(N)$ is $N(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})$. Suppose that $Np(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})$ of them are expected to be matched in the optimal allocation. Here $p$ depends on $b_{1},b_{2},r$.
If $\mu_{b_{1},b_{2},r,1}/\mu_{b_{1},b_{2},r,0}>\overline{\mu}_{r,1}/\overline{\mu}_{r,0}$, we can only match $np\mu_{b_{1},b_{2},r,0}$ vertices of type $(b_{1},b_{2},r,0)$ and $np\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}/\overline{\mu}_{r,0}$ vertices of type $(b_{1},b_{2},r,1)$ to balance the selection rates of two subgroups. The local relative efficiency loss is no greater than
$1-\frac{np\mu_{b_{1},b_{2},r,0}+np\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}/\overline{\mu}_{r,0}}{np(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})}=\frac{\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}-\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,0}}.$
If $\mu_{b_{1},b_{2},r,1}/\mu_{b_{1},b_{2},r,0}\leq\mu_{r,1}/\mu_{r,0}$, we can only match $np\mu_{b_{1},b_{2},r,1}$ vertices of type $(b_{1},b_{2},r,1)$ and $np\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}/\overline{\mu}_{r,1}$ vertices of type $(b_{1},b_{2},r,0)$ to balance the selection rates of two subgroups. The local relative efficiency loss is no greater than
$1-\frac{np\mu_{b_{1},b_{2},r,1}+np\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}/\overline{\mu}_{r,1}}{np(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})}=\frac{\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}-\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,1}}.$
Therefore, the overall relative efficiency loss is no greater than
almost surely as $n\to\infty$. ∎
Proof of Proposition 4.
Following the notation in Ashlagi and Roth (2011), an X-Y pair has a donor of blood type Y and a patient of blood type X. Without loss of generality, we assume there are more A-B pairs than B-A pairs. By Proposition 5.2 in Ashlagi and Roth (2011), almost surely as $n\to\infty$, there is an optimal allocation such that
  1. (1)
    every pair X-X is matched in a 2-way or a 3-way exchange with other X-X pairs, for X=O,A,B,AB;
  2. (2)
    every B-A pair is matched in a 2-way exchange with A-B pairs;
  3. (2)
    every AB-B pair is matched in a 2-way exchange with B-AB pairs;
  4. (3)
    every AB-A pair is matched in a 2-way exchange with A-AB pairs;
  5. (4)
    every AB-O pair is matched in a 3-way exchange with A-AB pairs and O-A pairs;
  6. (5)
    every A-O pair is matched in a 2-way exchange with O-A pairs
  7. (6)
    every B-O pair is either matched in a 2-way exchange with O-B pairs or in a 3-way exchange with A-B pairs, which are not matched with B-A pairs, and O-A pairs.
From the above optimal allocation, all the X-X, B-A, A-B, AB-B, AB-A, AB-O, A-O, B-O pairs are fully matched, the B-AB, A-AB, O-A, O-B pairs are partially matched, and no O-AB pairs are matched. By Proposition 2, within the blood type pairs that are partially matched, we can give priority to any specific sensitive attribute level $a$ among patients with sensitization level $r$. Based on such an explicit optimal allocation rule, we give an upper bound for the asymptotic price of fairness.
Let $\rho^{-1}$ be the probability that a random patient and a random donor are incompatible. Let $q_{j}$ be the frequency probability of sensitization level $R=r_{j}$. According to the optimal allocation rule, all the X-X, B-A, A-B, AB-B, AB-A, AB-O, A-O, B-O pairs are fully matched, the B-AB, A-AB, O-A, O-B pairs are partially matched, and no O-AB pairs are matched. We partition $V_{i,j}=\{V_{i,j,full}\cup V_{i,j,partial}\cup V_{i,j,none}\}$, where $V_{i,j,full}$ denotes the set of fully matched blood type pairs, $V_{i,r,partial}$ denotes the set of partially matched blood type pairs, and $V_{i,j,none}$ denotes the set of O-AB pairs, within $V_{i,j}$, the set of incompatible pairs of sensitive attribute level $a$ and sensitization level $r$. By Proposition 2, we can arbitrarily arrange the matched pairs in $V_{i,j,partial}$ to achieve fairness. The expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is equal to $q_{j}$ times the expected proportion of matched pairs in $\cup_{j}\{V_{1,j,partial}\cup V_{0,j,partial}\}$, i.e.,
$\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}\left\{\mu_{B\mid k}\mu_{AB\mid k}+(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k}+(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k}+\mu_{O\mid k}\mu_{B\mid k}\right\}.$
Since the expected proportion of $V_{i,j,full}$ is
$\rho\mu_{i}q_{j}\{r_{j}(\mu_{O\mid i}^{2}+\mu_{A\mid i}^{2}+\mu_{B\mid i}^{2}+\mu_{AB\mid i}+\mu_{A\mid i}\mu_{O\mid i}+\mu_{B\mid i}\mu_{O\mid i})+2\mu_{A\mid i}\mu_{B\mid i}\}=\rho q_{j}T_{i}(r_{j})$
and the expected proportion of $V_{i,r}$ is
$\rho q_{j}T_{i}(r_{j})+\rho\mu_{i}q_{j}(\mu_{B\mid i}\mu_{AB\mid i}+\mu_{A\mid i}\mu_{AB\mid i}+\mu_{O\mid i}\mu_{A\mid i}+\mu_{O\mid i}\mu_{B\mid i}+\mu_{O\mid i}\mu_{AB\mid i})=\rho q_{j}S_{i}(r_{j})$
the relative proportion of fully matched pairs within $V_{i,j}$ is $T_{i}(r_{j})/S_{i}(r_{j})$.
We first consider the case when $T_{1}(r_{j})/S_{1}(r_{j})<T_{0}(r_{j})/S_{0}(r_{j})$, under which we should give priority to sensitive attribute level $A=1$ within the sensitization level $R=r$ to balance the selection rates of subgroups.
  1. 1.
    For B-AB pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}\mu_{B\mid k}\mu_{AB\mid k}$
    while the expected proportion of B-AB pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{B\mid 1}\mu_{AB\mid 1}$. Thus, we can arrange the matched B-AB pairs in $V_{1,j,partial}$ and $V_{0,j,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{B\mid k}\mu_{AB\mid k},2\mu_{1}\mu_{B\mid 1}\mu_{AB\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{B\mid k}\mu_{AB\mid k}\right\}.$
  2. 2.
    For A-AB pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k}$
    while the expected proportion of A-AB pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{A\mid 1}\mu_{AB\mid 1}$. Thus, we can arrange the matched A-AB pairs in $V_{1,r,partial}$ and $V_{0,r,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k},2\mu_{1}\mu_{A\mid 1}\mu_{AB\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k}\right\}.$
  3. 3.
    For O-A pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k}$
    while the expected proportion of O-A pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{O\mid 1}\mu_{A\mid 1}$. Thus, we can arrange the matched O-A pairs in $V_{1,r,partial}$ and $V_{0,r,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k},2\mu_{1}\mu_{O\mid 1}\mu_{A\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k}\right\}.$
  4. 4.
    For O-B pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}\mu_{O\mid k}\mu_{B\mid k}$
    while the expected proportion of O-B pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{1,B}\mu_{1,AB}$. Thus, we can arrange the matched O-B pairs in $V_{1,r,partial}$ and $V_{0,r,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{O\mid k}\mu_{B\mid k},2\mu_{1}\mu_{O\mid 1}\mu_{B\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{O\mid k}\mu_{B\mid k}\right\}.$
The above process leads to a total difference in expected proportion equal to $\rho q_{j}R_{1}$. If $(T_{1}(r_{j})+R_{1})/S_{1}(r_{j})\geq T_{0}(r_{j})/S_{0}(r_{j})$, by continuity, we can always arrange the matched pairs such that the fairness can be perfectly achieved without losing any efficiency. That is, the price of fairness is 00. However, if $(T_{1}(r_{j})+R_{1})/S_{1}(r_{j})<T_{0}(r_{j})/S_{0}(r_{j})$, we can choose to drop some matched pairs in $V_{0,j,full}$ to sacrifice efficiency for fairness. We solve $x$ from the following equation
$\frac{T_{1}(r_{j})+R_{1}}{S_{1}(r_{j})}=\frac{T_{0}(r_{j})-x}{S_{0}(r_{j})}$
and the solution is
$x=\frac{T_{0}(r_{j})S_{1}(r_{j})-T_{1}(r_{j})S_{0}(r_{j})-R_{1}S_{0}(r_{j})}{S_{1}(r_{j})}.$
Thus, the price of fairness within the stratum $R=r_{j}$ is no greater than $x/U(r_{j})$.
We next consider the case when $T_{1}(r_{j})/S_{1}(r_{j})\geq T_{0}(r_{j})/S_{0}(r_{j})$. Similarly, the price of fairness within the level $R=r_{j}$ is no greater than
$\max\left\{\frac{S_{1}T_{0}-S_{0}T_{1}-S_{0}R_{1}}{S_{1}U},\frac{S_{0}T_{1}-S_{1}T_{0}-S_{1}R_{0}}{S_{0}U},0\right\}.$
Therefore, among the whole population, the price of fairness is no greater than
$\max_{r\in\{r_{1},...,r_{M}\}}\max\left\{\frac{S_{1}T_{0}-S_{0}T_{1}-S_{0}R_{1}}{S_{1}U},\frac{S_{0}T_{1}-S_{1}T_{0}-S_{1}R_{0}}{S_{0}U},0\right\}.$
@@ Appendix B More simulations with failure-aware strategies
We repeat the data generating process in Sections 5.1 and 5.2, respectively. With additional vertex and edge uncertainties, let the failure probability $p_{v}$ and $p_{v_{i},v_{j}}$ be independently sampled from a uniform distribution $U(0,0.3)$. We consider the subset-recourse strategy and choose S to be the set of relevant subsets of size $(3,1)$.
Figures S1 and S2 present the average selection rates of different fairness criteria within each subgroup. We do not display the results under weaker fairness constraints of group fairness, individual fairness, and our novel fairness, because they are very close to the results without any fairness constraint. The numerical performance with failure-aware strategies is similar to that without failure-aware strategies.

### v2

@@ Appendix A Proofs
We work in the random graph model of Section 4 specialized to the case where the objective is to maximize the number of transplants (no recourse strategies), and we adopt the additional structural assumptions stated before Proposition 3.
Notational convention. In the rest of the paper we index blood types of a pair as
$(b_{1},b_{2})=(\text{donor blood type},\ \text{patient blood type}).$
By contrast, Ashlagi and Roth describe a pair of type XXYY when the patient has blood type $X$ and the donor has blood type $Y$. Thus their type XXYY corresponds to our ordered pair
$(b_{1},b_{2})=(Y,X).$
Whenever we refer below to XXYY pair types or to their classification as self-/over-/under-demanded from Ashlagi and Roth (2014), this is always in the patient–donor sense of Ashlagi–Roth; when we translate those statements into our formulas in terms of $\mu_{b_{1}\mid a}$ and $\mu_{b_{2}\mid a}$, we use the above correspondence.
The proof has four main steps.
  • Step 1 recalls the blood-type structure of an asymptotically optimal allocation without fairness constraints, following Ashlagi and Roth (2014), and translates it into our $(b_{1},b_{2})$ notation.
  • Step 2 quantifies how much matched “mass” can be reallocated between $A=0$ and $A=1$ within partially matched blood-type classes; this yields $R_{0}$ and $R_{1}$.
  • Step 3 uses this flexibility and, if necessary, discards some matches to enforce equality of selection probabilities across $A=0$ and $A=1$ at each sensitization level $r$, and computes the resulting loss.
  • Step 4 uses Proposition 1 to argue that this type-level construction can be implemented in the random graphs $G_{N}$ up to $o(N)$ loss and that the weaker fairness constraint (4) has no higher asymptotic price.
Step 1: Structure of the unconstrained optimal allocation. Fix $N$ and let $G_{N}$ be the random compatibility graph on $N$ incompatible pairs. For each $a\in\{0,1\}$ and $r\in\{r_{1},\dots,r_{M}\}$, write
$V_{a,r}:=\{v\in V_{N}:A(v)=a,\ R(v)=r\}.$
By the strong law and the multiplicative structure,
$\frac{|V_{a,r}|}{N}\;\longrightarrow\;\sum_{b_{1},b_{2}}\mu_{b_{1},b_{2},r,a}=c\,r\,\mu_{a}\quad\text{a.s.\ as }N\to\infty.$
In Ashlagi and Roth (2014), pair types are indexed as XXYY with $X$ the patient blood type and $Y$ the donor blood type. Translating into our $(b_{1},b_{2})$ notation via $(b_{1},b_{2})=(Y,X)$, a type XXYY in Ashlagi and Roth (2014) corresponds to patient type $X$ and donor type $Y$, i.e. our $(b_{1},b_{2})=(Y,X)$.
Using that convention, Ashlagi and Roth (2014) partition the $4\times 4$ blood-type pairs XXYY as follows:
  • self-demanded: XXXX (patient and donor have the same blood type);
  • over-demanded: XXYY with $Y\to X$ and $X\neq Y$ (donor offers a more flexible blood type than the patient needs);
  • under-demanded: XXYY with $X\to Y$ and $X\neq Y$ (patient seeks a more flexible type than the donor offers);
  • reciprocally demanded: AABB and BBAA.
Under the numerical assumptions on blood-type frequencies and sensitization, they show that in almost every large pool there exists an optimal allocation (with exchanges of length at most $3$) whose limiting structure is deterministic in terms of these blood-type classes. After translating their XXYY types into our $(b_{1},b_{2})$ notation, the key consequences within each group $(A=a,R=r)$ are:
  • all self-demanded types are asymptotically fully matched;
  • all AABB and BBAA patient–donor types are asymptotically fully matched via 2-cycles and 3-cycles;
  • all AB–O patient–donor types are matched using 3-cycles involving overdemanded types;
  • all remaining overdemanded types are matched to their corresponding underdemanded types in 2-cycles;
  • only the genuinely underdemanded types beyond available overdemanded partners remain unmatched.
Counting these matched classes in our $(b_{1},b_{2})$ notation, and using the product form
$\mu_{b_{1},b_{2},r,a}=c\,r\,\mu_{a}\,\mu_{b_{1}\mid a}\mu_{b_{2}\mid a},$
gives exactly the expressions $T_{a}(r)$ and $S_{a}(r)$ stated before the proposition. In particular, for each $a$ and $r$ the unconstrained selection probability in group $(A=a,R=r)$ converges to
$p_{a}(r):=\frac{T_{a}(r)}{S_{a}(r)}.$
Aggregating across $a$ gives the total matchable mass at level $r$, proportional to $Q(r)$ as defined in the statement.
Step 2: Reallocatable mass between groups and $R_{a}$. The Ashlagi–Roth construction leaves some blood-type classes (particularly those involving underdemanded types) only partially utilized. Within such classes we can reassign which fraction of matched pairs comes from $A=0$ versus $A=1$, provided we respect:
  • for each blood-type class, the total matched mass in that class; and
  • for each group $A=a$, the total number of available pairs of that class.
For each “bottleneck” class (e.g. patient–donor BB–AB, OO–AB, AA–AB, AA–O, AB–O, OO–B in the Ashlagi–Roth notation) this gives a small linear program in two variables (mass from $A=0$ and from $A=1$). Solving these programs and translating back to our $(b_{1},b_{2})$ representation gives the quantities $R_{a}$ as in the proposition:
$R_{a}=$ $\min\left\{\phi_{B,AB},2\mu_{a}\mu_{B\mid a}\mu_{AB\mid a}-\phi_{B,AB}\right\}+\min\left\{\phi_{O,AB}+\phi_{A,AB},2\mu_{a}\mu_{A\mid a}\mu_{AB\mid a}-\phi_{O,AB}-\phi_{A,AB}\right\}$
$+\min\left\{\phi_{A,O}+\phi_{AB,O},2\mu_{a}\mu_{O\mid a}\mu_{A\mid a}-\phi_{A,O}-\phi_{AB,O}\right\}+\min\left\{\phi_{O,B},2\mu_{a}\mu_{O\mid a}\mu_{B\mid a}-\phi_{O,B}\right\}.$
Informally, $R_{a}$ is the maximal amount of matched mass that can be shifted in favour of group $A=a$ by reassigning responsibility for matches within these partially utilized blood-type classes.
Step 3: Enforcing equality of selection probabilities at each $r$. Fix a sensitization level $r$. We aim for the stronger condition
# $\frac{\text{\# matches in }(A=1,R=r)}{|V_{1,r}|}=\frac{\text{\# matches in }(A=0,R=r)}{|V_{0,r}|},$
i.e. $p_{1}(r)$ and $p_{0}(r)$ are made equal. This is stricter than (4), so any loss bound obtained here is valid for our weaker fairness constraint.
Suppose first that $p_{1}(r)<p_{0}(r)$, so group $A=1$ is disadvantaged.
(a) Using $R_{1}$ to help $A=1$. By reallocating within partially matched classes we can increase the matched mass of $(A=1,R=r)$ by at most $R_{1}$, without changing the total mass at level $r$. This yields a new numerator $T_{1}(r)+R_{1}$ for $A=1$ and reduces the numerator for $A=0$ by the same $R_{1}$.
If this already suffices for equality,
$\frac{T_{1}(r)+R_{1}}{S_{1}(r)}\;\geq\;\frac{T_{0}(r)}{S_{0}(r)},$
then equality can be achieved with no loss of total matches at level $r$.
(b) If reallocation is insufficient, discarding from $A=0$. Otherwise we must additionally discard matches from $(A=0,R=r)$. Let $x(r)\geq 0$ be the extra amount of matched mass we remove from group $A=0$ at level $r$, beyond the $R_{1}$ already reallocated. Equality of selection probabilities requires
$\frac{T_{1}(r)+R_{1}}{S_{1}(r)}\;=\;\frac{T_{0}(r)-x(r)}{S_{0}(r)},$
hence
$x(r)=\frac{S_{0}(r)T_{1}(r)-S_{1}(r)T_{0}(r)-S_{0}(r)R_{1}}{S_{1}(r)}.$
(c) Symmetric case $p_{1}(r)\geq p_{0}(r)$. If $p_{1}(r)\geq p_{0}(r)$, the roles of the two groups are reversed: we can first reallocate up to $R_{0}$ in favour of $A=0$ and, if needed, discard an additional amount $x^{\prime}(r)\geq 0$ from group $A=1$, where
$x^{\prime}(r)=\frac{S_{1}(r)T_{0}(r)-S_{0}(r)T_{1}(r)-S_{1}(r)R_{0}}{S_{0}(r)}.$
(d) Relative loss at level $r$. At level $r$, the total matchable mass in the unconstrained optimum is proportional to $Q(r)$; hence the relative loss (price of fairness) at level $r$ is at most
$\max\left\{\frac{S_{1}(r)T_{0}(r)-S_{0}(r)T_{1}(r)-S_{0}(r)R_{1}}{S_{1}(r)Q(r)},\;\frac{S_{0}(r)T_{1}(r)-S_{1}(r)T_{0}(r)-S_{1}(r)R_{0}}{S_{0}(r)Q(r)},\;0\right\},$
which is exactly the quantity appearing in the statement.
Step 4: Implementability via subgroup prioritization and Proposition 1. Up to now the construction has been at the level of limiting type proportions. We now argue that it can be implemented in the random graphs $G_{N}$ (with exchanges of bounded length) up to $o(N)$ loss.
For each triple $(b_{1},b_{2},r)$ and $a\in\{0,1\}$, the reallocation and discarding operations in Step 3 determine whether, at that type, the fair allocation effectively prioritizes group $A=1$ or $A=0$. Collect these choices in a priority set $\mathcal{P}$ as in Section 4.1, and consider the optimization problem with subgroup-prioritization constraints (6)–(7) associated with $\mathcal{P}$.
By Proposition 1, there exists a sequence $\varepsilon_{N}\downarrow 0$ such that, almost surely, we can impose the asymptotic priority requirements (7) for all $(b_{1},b_{2},i,j)\in\mathcal{P}$ while losing only $o(N)$ matches relative to the unconstrained optimum. In other words, the prioritized solutions implement, up to $o(N)$, the type-level pattern constructed in Step 3.
Within this prioritized class, we can then carry out the same trimming as in Step 3 to enforce equality of selection probabilities at each $r$. The additional loss in the number of matches is, at level $r$, bounded by the relative loss above, and hence the overall asymptotic price of fairness is bounded by
$\max_{r\in\{r_{1},\dots,r_{M}\}}\max\left\{\frac{S_{1}(r)T_{0}(r)-S_{0}(r)T_{1}(r)-S_{0}(r)R_{1}}{S_{1}(r)Q(r)},\;\frac{S_{0}(r)T_{1}(r)-S_{1}(r)T_{0}(r)-S_{1}(r)R_{0}}{S_{0}(r)Q(r)},\;0\right\}.$
Finally, the original fairness constraint (4) requires only that the difference in selection probabilities at each $r_{j}$ be bounded by $l_{j}$, not that they be exactly equal. Thus an optimal solution under (4) is at least as efficient as the equalized allocations we have constructed, so its asymptotic price of fairness is no larger than the bound stated in Proposition 3.
This completes the proof. ∎
@@ Appendix B More simulations with failure-aware strategies
We consider a similar data-generating process in Section 5.1 but with $N=50$, and we repeat exactly the same data-generating process in Section 5.2. With additional vertex and edge uncertainties, let the failure probability $p_{v}$ and $p_{v_{i},v_{j}}$ be independently sampled from a uniform distribution $U(0,0.3)$. We consider the subset-recourse strategy and choose $\mathcal{S}$ to be the set of relevant subsets of size $(3,1)$.
Figures S1 and S2 present the average selection rates of different fairness criteria within each subgroup. The numerical performance with failure-aware strategies is similar to that without failure-aware strategies.